# When Are Bias-Free ReLU Networks Effectively Linear Networks?

### v2

@@ 1 Introduction
By revealing regimes where bias-free ReLU networks behave like linear networks, we provide an accessible way of understanding ReLU networks within these regimes, as well as a cautionary note that studying nonlinear behaviors generally requires stepping beyond these regimes. This perspective leverages insights from linear networks, which are simpler and thus enjoy much richer theoretical results than ReLU networks (Baldi & Hornik, 1989; Fukumizu, 1998; Saxe et al., 2014; 2019; Arora et al., 2018; Ji & Telgarsky, 2019; Lampinen & Ganguli, 2019; Gidel et al., 2019; Tarmoun et al., 2021; Braun et al., 2022; Ziyin et al., 2022). For example, we are able to give closed-form time-course solutions to certain two-layer ReLU networks outside the lazy learning regime in Corollary 8. Our findings suggest that the bias terms in a ReLU network play an important role in learning nonlinear tasks. Our contributions are the following:
@@ 1.1 Related Work
A few other works have alluded to the connections between two-layer ReLU and linear networks. Sarussi et al. (2021) discovered that two-layer bias-free leaky ReLU networks converge to a decision boundary that is very close to linear when the teacher model is linear. Their theoretical results assume that the second layer is fixed while we train all layers of the network. Saxe et al. (2022) studied gated deep linear networks and found they closely approximate a two-layer bias-free ReLU network trained on an XOR task. But they did not generalize the connection between gated linear networks and ReLU networks beyond the XOR case. Boursier & Flammarion (2024) gave an example dataset with three scalar input data points, in which two-layer bias-free (leaky) ReLU networks converge to the linear, ordinary least square estimator. Holzmüller & Steinwart (2022) studied two-layer leaky ReLU networks with bias and found that they perform linear regression on certain data distributions with scalar input, because the bias fails to move far away from their initialization at zero.
@@ 2.1 Two-Layer Bias-Free (Leaky) ReLU and Linear Networks
$f({\bm{x}};{\bm{W}})={\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}})=\sum_{h=1}^{H}w_{2h}\sigma({\bm{w}}_{1h}{\bm{x}}),\quad\text{where }\sigma(z)=\max(z,\alpha z),\alpha\in[0,1].$ (1)
@@
where ${\bm{\Sigma}}$ denotes the input data covariance and ${\bm{\beta}}$ denotes the input-output correlation,
@@ 3 Network Expressivity
We first examine the expressivity of bias-free ReLU networks. It is well known that standard ReLU networks are universal approximators (Hornik et al., 1989; Pinkus, 1999) while bias-free ReLU networks are not since they can only express positively homogeneous functions, i.e., $g(a{\bm{x}})=ag({\bm{x}})\,\forall a>0$. Moreover, Section 3.1 shows that two-layer bias-free ReLU networks cannot express any odd function except linear functions. Section 3.2 shows that deep bias-free ReLU networks are more expressive than two-layer ones, but are still limited to positively homogeneous functions.
@@ 3.1 Two-Layer Bias-Free (Leaky) ReLU Networks
Two-layer bias-free (leaky) ReLU networks can only express a linear function plus a positively homogeneous even function.
@@
An arbitrary two-layer (leaky) ReLU network can be written as
$\sum_{h=1}^{H}w_{2h}\sigma({\bm{w}}_{1h}{\bm{x}})=\sum_{h=1}^{H}w_{2h}\left[\frac{1+\alpha}{2}{\bm{w}}_{1h}{\bm{x}}+\frac{1-\alpha}{2}|{\bm{w}}_{1h}{\bm{x}}|\right],\vspace{-5ex}$ (8)
@@ 4.1 Symmetric Datasets
Section 3.1 has proven that two-layer bias-free ReLU networks cannot express odd functions except linear functions. We now show that under the Condition 3 on the dataset, two-layer bias-free ReLU networks not only find a linear solution but also have the same learning dynamics as a two-layer linear network.
@@
For infinite data, the first part of Condition 3 includes common distributions, such as any Gaussian distribution with zero mean. For finite data, the first part of Condition 3 means that if ${\bm{x}}$ is present in the dataset, $-{\bm{x}}$ is also present. Condition 3 includes the dataset studied in Lyu et al. (2021). They considered linearly separable binary classification tasks with a data augmentation procedure in which $(-{\bm{x}},-y)$ is added to the dataset if $({\bm{x}},y)$ is in the dataset. We have the same assumption on the input data distribution but relax the assumption on the target model from being linearly separable to being odd.
@@
At initialization, there exist an unit vector ${\bm{r}}$ such that ${\bm{W}}_{1}={\bm{W}}_{2}^{\top}{\bm{r}}^{\top}$ and the second-layer weight ${\bm{W}}_{2}$ has equal L2 norms for its positive and negative elements.
Remark 6.
Under Condition 3, the dynamics of two-layer bias-free (leaky) ReLU networks and linear networks initialized with small weights can both be approximated by a linear differential equation in the early phase of learning. Thus, the weights of the ReLU network and the weights of the linear network form the same rank-one structure and the left singular vector of ${\bm{W}}_{1}$ is aligned with ${\bm{W}}_{2}$, as proven in Theorem 15. At the end of the early phase, the weights are rank-one and aligned with bounded errors. To simplify the analysis, Assumption 5 assumes the weights are exactly rank-one and aligned, which is justified by the initialization approaching 00. We also assume, for simplicity, that the second-layer weights have equal L2 norm for their positive and negative elements, which is justified by the width approaching infinity. Simulations in LABEL:fig:L2-sym-error show that errors are small in the finite case as well.
Under Assumptions 5 and 3, the learning dynamics of two-layer bias-free (leaky) ReLU networks reduces to (see Section C.2)
@@
Except for the constant coefficients, these equations are the same as the ones for the linear network given in Equation 3. Thus, apart from the fact that learning is $(\alpha+1)/2$ times slower and the weights are $\sqrt{2/(\alpha+1)}$ times larger, the ReLU network has the same learning dynamics as its linear counterpart. We formally state this equivalence below.
Figure 2: Two-layer bias-free (leaky) ReLU networks can evolve like a linear network. (a) Loss curves with different leaky ReLU parameter $\alpha$ (note $\alpha=1$ is a linear network). The simulations match the theoretical solutions in Equation 13. The loss converges to global minimum, which is not zero due to the restricted expressivity of two-layer bias-free ReLU networks. (b) The simulated loss curves are plotted against a rescaled time axis; they collapse to one curve, demonstrating the networks are implementing the same linear function as in Equation 11. The error, defined as $\left\|\sqrt{\frac{\alpha+1}{2}}{\bm{W}}\left(\frac{2}{\alpha+1}t\right)-{\bm{W}}^{\mathrm{lin}}(t)\right\|/\left\|{\bm{W}}^{\mathrm{lin}}(t)\right\|$, is less than $0.3\%$, demonstrating that the weights in the (leaky) ReLU network are close to the weights in the linear network as in Equation 12. The errors are not exactly zero because the initialization is small but nonzero in the simulations. Experimental details are provided in Appendix H.
Theorem 7.
A two-layer (leaky) ReLU network and a linear network are trained with square or logistic loss starting from weights which differ by a scale factor, ${\bm{W}}(0)=\sqrt{2/(\alpha+1)}\,{\bm{W}}^{\mathrm{lin}}(0)$. Under Condition 3 on the dataset and Assumption 5 on the initial weights, we have that $\forall\,t\geq 0$, Assumption 5 remains valid and:
@@
As a sanity check, setting $\alpha=1$ yields a linear network, where ${\bm{W}}$ is identical to ${\bm{W}}^{\mathrm{lin}}$. As an example, setting $\alpha=0$ yields a ReLU network, where learning slows down by half, explaining a previous empirical observation (Zhang et al., 2024). We validate Theorem 7 and the plausibility of Assumption 5 with numerical simulations in Figure 2. Additionally, we provide theoretical proof that Theorem 7 holds with L2 regularization and empirical evidence that some of Theorem 7 hold with large initialization and a moderately large learning rate in Sections C.4, C.5 and C.6.
@@
Under the same conditions as Theorem 7, the two-layer bias-free (leaky) ReLU network converges to a linear solution $f({\bm{x}};{\bm{W}}(\infty))={{\bm{w}}^{*}}^{\top}{\bm{x}}$. For square loss, ${\bm{w}}^{*}$ is the ordinary least squares solution, ${\bm{\Sigma}}^{-1}{\bm{\beta}}$, which is the global minimum. For linearly separable binary classification with logistic loss, ${\bm{w}}^{*}$ is the max-margin solution.
@@ 4.2 Orthogonal and XOR Datasets
In Section 4.1 we considered symmetric datasets where a two-layer bias-free ReLU network evolves like one linear network. We now present two datasets where a two-layer bias-free ReLU network evolves like multiple independent linear networks.
@@
The first dataset has orthogonal inputs, which is a common setting studied by prior literature (Boursier et al., 2022; Telgarsky, 2023; Frei et al., 2023b; c; Kou et al., 2023). In particular, we use the same dataset as Boursier et al. (2022, Figure 3), which consists of two orthogonal data points, as shown in LABEL:fig:ortho-dataset. We train a two-layer bias-free ReLU network on this dataset to reproduce the loss curve in (Boursier et al., 2022, Figure 3). We then train two two-layer linear networks on each data point separately. We find that the timing and the amount of the loss drop overlap with the loss curves of the two linear networks as shown in LABEL:fig:ortho-loss. To understand this overlap, we plot the first-layer weights of the ReLU network in black arrows in LABEL:fig:ortho-dataset and find that the weights align with either one of the two data points. Since the two directions are orthogonal, the learning dynamics of the two groups of neurons decouple, as derived in Appendix D. Each group of neurons evolves like a linear network trained on that single data point. Hence, weights in the ReLU network evolve like a linear network trained on either one of the two data points separately. The same applies to learning with logistic loss, as shown in Figure 7, which was not covered in Boursier et al. (2022).
@@
Figure 3: Two-layer bias-free ReLU networks can evolve like several linear networks. (a) An orthogonal input dataset used in (Boursier et al., 2022, Figure 3). The $+$ and -- signs represent data points with $+1$ and $-1$ labels respectively. Their different colors are used only to distinguish the loss curves. The black arrows are the first-layer weights at convergence. (b) The loss curve of the ReLU network overlaps with two linear networks trained on each of the two data points respectively. (c) An XOR-like dataset. (d) The loss curve of the ReLU network overlaps with four linear networks trained on each of the four data points separately. Details: We use summed (instead of averaged) square loss for this figure. The initial losses are vertically aligned to help illustrate the overlap. More details are in Appendix H.
@@
We observe similar behavior in the XOR-like task shown in LABEL:fig:xor-dataset, which is neither linearly separable nor odd. XOR-like datasets are also a common setting in prior literature (Saxe et al., 2022; Frei et al., 2023a; Meng et al., 2024; Xu et al., 2024; Glasgow, 2024). As shown in LABEL:fig:xor-loss, we find that the loss curves of a two-layer bias-free ReLU network trained on the XOR task overlap with four linear networks trained on each data point separately. In this case, the dynamics of multiple linear networks well approximate that of a ReLU network, even though the ReLU network learns a nonlinear function.
@@
In LABEL:fig:ortho-loss and LABEL:fig:xor-loss, the loss curves go through multiple decreases, each corresponding to learning a data point. Similar behaviors were examined by Boursier et al. (2022); Xu et al. (2024) and characterized as saddle-to-saddle dynamics. The connections we find between ReLU and linear networks may help understand these behaviors in ReLU networks because saddle-to-saddle dynamics has been well studied for linear networks (Saxe et al., 2014; 2019; Jacot et al., 2021; Pesme & Flammarion, 2023).

### v3

@@ 1 Introduction
By revealing regimes where bias-free ReLU networks behave like linear networks, we provide an accessible way of understanding ReLU networks within these regimes, as well as a cautionary note that studying nonlinear behaviors generally requires stepping beyond these regimes. This understanding leverages insights from linear networks, which are simpler and thus enjoy much richer theoretical results than ReLU networks (Baldi & Hornik, 1989; Fukumizu, 1998; Saxe et al., 2014; 2019; Arora et al., 2018; Ji & Telgarsky, 2019; Lampinen & Ganguli, 2019; Gidel et al., 2019; Tarmoun et al., 2021; Braun et al., 2022; Ziyin et al., 2022). For example, we are able to give closed-form time-course solutions to certain two-layer ReLU networks outside the lazy learning regime in Corollary 9. Our findings suggest that the bias terms in a ReLU network play an important role in learning nonlinear tasks. Our contributions are the following:
@@ 1.1 Related Work
A few other works have alluded to the connections between two-layer ReLU and linear networks. Sarussi et al. (2021) discovered that two-layer bias-free leaky ReLU networks converge to a decision boundary that is very close to linear when the teacher model is linear. Their theoretical results assume that the second layer is fixed while we train all layers of the network. Saxe et al. (2022) studied gated deep linear networks and found they closely approximate a two-layer bias-free ReLU network trained on an XOR task. But they did not generalize the connection between gated linear networks and ReLU networks beyond the XOR case. Boursier & Flammarion (2024a) gave an example dataset with three scalar input data points, in which two-layer bias-free (leaky) ReLU networks converge to the linear, ordinary least square estimator. Holzmüller & Steinwart (2022) studied two-layer leaky ReLU networks with bias and found that they perform linear regression on certain data distributions with scalar input, because the bias fails to move far away from their initialization at zero.
@@ 2.1 Two-Layer Bias-Free (Leaky) ReLU and Linear Networks
$f({\bm{x}};{\bm{W}})={\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{x}})=\sum_{h=1}^{H}w_{2h}\sigma\left({\bm{w}}_{1h}^{\top}{\bm{x}}\right),\quad\text{where }\sigma(z)=\max(z,\alpha z),\alpha\in[0,1].$ (1)
@@
where ${\bm{\Sigma}}$ denotes the input covariance and ${\bm{\beta}}$ denotes the input-output correlation,
@@ 3 Network Expressivity
We first examine the expressivity of bias-free ReLU networks. It is well known that standard ReLU networks with bias are universal approximators (Hornik et al., 1989; Pinkus, 1999) while bias-free ReLU networks are not since they can only express positively homogeneous functions, i.e., $g(a{\bm{x}})=ag({\bm{x}})\,\forall a>0$. Moreover, Section 3.1 shows that two-layer bias-free ReLU networks cannot express any odd function except linear functions. Section 3.2 shows that deep bias-free ReLU networks are more expressive than two-layer ones, but are still limited to positively homogeneous functions.
@@ 3.1 Two-Layer Bias-Free (Leaky) ReLU Networks
The set of functions that can be expressed by two-layer bias-free (leaky) ReLU networks is a subset of the set of functions of the form: $f({\bm{x}})=h({\bm{x}})+g({\bm{x}})$, where $h({\bm{x}})$ is linear and $g({\bm{x}})$ is a positively homogeneous even function, meaning $g({\bm{x}})=g(-{\bm{x}})$ and $g(a{\bm{x}})=ag({\bm{x}})\,\forall a>0$.
@@
Notice that the (leaky) ReLU activation function admits a decomposition [^1] Decomposing the ReLU activation function into a linear and an even function is technically uncomplicated and has appeared in prior literature before, e.g., Ghorbani et al. (2021, Section 1.3) and Martinelli et al. (2024, page 4).: $\sigma(z)=\frac{1+\alpha}{2}z+\frac{1-\alpha}{2}|z|$. Thus, any two-layer (leaky) ReLU network can be written as
$f({\bm{x}};{\bm{W}})=\sum_{h=1}^{H}w_{2h}\sigma({\bm{w}}_{1h}{\bm{x}})=\sum_{h=1}^{H}w_{2h}\left[\frac{1+\alpha}{2}{\bm{w}}_{1h}{\bm{x}}+\frac{1-\alpha}{2}|{\bm{w}}_{1h}{\bm{x}}|\right],$ (8)
@@ 4.1 Symmetric Datasets
Section 3.1 has proven that two-layer bias-free ReLU networks cannot express odd functions except linear functions. We now show that under Condition 3 on the dataset, two-layer bias-free ReLU networks not only find a linear solution but also have the same learning dynamics as a two-layer linear network.
@@
For infinite data, the first part of Condition 3 includes common distributions, such as any Gaussian distribution with zero mean. For finite data, the first part of Condition 3 means that if ${\bm{x}}$ is present in the dataset, $-{\bm{x}}$ is also present. Condition 3 includes the dataset studied in Lyu et al. (2021). They considered linearly separable binary classification tasks with a data augmentation procedure in which $(-{\bm{x}},-y)$ is added to the dataset if $({\bm{x}},y)$ is in the dataset. We have the same assumption on the input data distribution but relax the assumption on the target function from being linearly separable to being odd.
The key implication of Condition 3 is that the input covariance matrix and the input-output correlation averaged over any half space are equal to those averaged over the entire space. We state this in Lemma 5 and prove it in Section B.2.
Lemma 5.
Let set ${\mathbb{S}}^{+}$ be an arbitrary half space divided by a hyperplane with normal vector ${\bm{r}}$, namely ${\mathbb{S}}^{+}=\{{\bm{x}}\in{\mathbb{R}}^{D}|{\bm{r}}^{\top}{\bm{x}}>0\}$. Under Condition 3, we have $\forall{\bm{r}}$
$\left\langle{\bm{x}}{\bm{x}}^{\top}\right\rangle_{{\mathbb{S}}^{+}}={\bm{\Sigma}},\quad\left\langle{\bm{x}}y({\bm{x}})\right\rangle_{{\mathbb{S}}^{+}}={\bm{\beta}}.$ (10)
Under Condition 3, two-layer bias-free (leaky) ReLU networks initialized with small random weights evolve approximately according to a linear differential equation in the early phase of learning. We can solve the approximate linear differential equation and bound the errors of the approximation, leading to the following lemma.
Lemma 6.
We define the initialization scale as $w_{\mathrm{init}}=\max(\|{\bm{W}}_{1}(0)\|,\|{\bm{W}}_{2}(0)\|)$. For time $t<\frac{\tau}{s+\operatorname{Tr}{\bm{\Sigma}}}\ln\frac{1}{w_{\mathrm{init}}}$, the solution to the dynamics of two-layer (leaky) ReLU networks starting from small initialization exhibits exponential growth along one direction with small errors
The proof for Lemma 6 can be found in Section C.1. Lemma 6 indicates that the weights of the ReLU network and the linear network form the same rank-one structure in the early phase, differing only in the speed of exponential growth. At the end of the early phase, the weights are aligned and rank-one with bounded errors, $O(w_{\mathrm{init}})$. For simplicity, we will assume in Assumption 7 that the weights are exactly aligned and rank-one, which is justified when the initialization is infinitesimal, $w_{\mathrm{init}}\to 0$. Assumption 7 further assumes that ${\bm{W}}_{2}$ has equal L2 norms for their positive and negative elements. This assumption is supported by the fact that ${\bm{W}}_{2}$ is proportional to ${\bm{r}}_{1}$, which follows a zero-mean Gaussian distribution because the initial weights ${\bm{W}}_{1}(0),{\bm{W}}_{2}(0)$ are sampled from a zero-mean Gaussian distribution. As the network width approaches infinity, ${\bm{W}}_{2}$ consists of infinitely many zero-mean Gaussian random samples, which have equal L2 norms for their positive and negative elements.
@@
At initialization, there exists an unit vector ${\bm{r}}$ such that ${\bm{W}}_{1}={\bm{W}}_{2}^{\top}{\bm{r}}^{\top}$, and the L2 norms of the positive and negative elements in ${\bm{W}}_{2}$ are equal, that is $\|\max({\bm{W}}_{2},0)\|=\|\max(-{\bm{W}}_{2},0)\|$ where $\max(\cdot)$ is applied element-wise.
Theorem 8.
A two-layer (leaky) ReLU network and a linear network are trained with square or logistic loss starting from weights which differ by a scale factor, ${\bm{W}}(0)=\sqrt{2/(\alpha+1)}\,{\bm{W}}^{\mathrm{lin}}(0)$. Under Condition 3 on the dataset and Assumption 7 on the initial weights, the learning dynamics of the two-layer (leaky) ReLU network reduces to
@@
For all $t\geq 0$, Assumption 7 remains valid and the following hold:
@@
The proof for Theorem 8 can be found in Section C.2. The key step is that the reduced dynamics of the ReLU network given in Equation 12 is the same as that for the linear network in Equation 3, modulo the constant coefficients. Hence, apart from the fact that learning is $(\alpha+1)/2$ times slower and the weights are $\sqrt{2/(\alpha+1)}$ times larger, the ReLU network has the same learning dynamics as its linear counterpart.
We validate Theorem 8 and the plausibility of Assumption 7 with numerical simulations in Figure 3. In LABEL:fig:L2-sym-error, the initialization is small random Gaussian weights and thus does not satisfy Assumption 7, yet Theorem 8 holds with small errors (less than $0.3\%$). Furthermore, we provide theoretical proof that Theorem 8 holds with L2 regularization and empirical evidence that some of Theorem 8 hold with large initialization and a moderately large learning rate in Sections C.4, C.5 and C.6.
Figure 3: Two-layer bias-free (leaky) ReLU networks can evolve like a linear network. (a) Loss curves with different leaky ReLU parameters $\alpha$ (note $\alpha=1$ is a linear network). The simulations match the theoretical solutions in Equation 15. The loss converges to global minimum, which is not zero due to the restricted expressivity of two-layer bias-free ReLU networks. (b) The simulated loss curves are plotted against a rescaled time axis; they collapse to one curve, demonstrating the (leaky) ReLU and linear networks are implementing the same linear function as in Equation 13. The error, defined as $\left\|\sqrt{\frac{\alpha+1}{2}}{\bm{W}}\left(\frac{2}{\alpha+1}t\right)-{\bm{W}}^{\mathrm{lin}}(t)\right\|/\left\|{\bm{W}}^{\mathrm{lin}}(t)\right\|$, is less than $0.3\%$, demonstrating that the weights in the (leaky) ReLU network are close to the weights in the linear network as in Equation 14. The errors are not exactly zero because the initial weights are sampled from a zero-mean Gaussian distribution, which does not satisfy Assumption 7 but better reflects practical initialization schemes. Experimental details are provided in Appendix H.
@@
Under the same conditions as Theorem 8, the two-layer bias-free (leaky) ReLU network converges to a linear solution $f({\bm{x}};{\bm{W}}(\infty))={{\bm{w}}^{*}}^{\top}{\bm{x}}$. For square loss, ${\bm{w}}^{*}$ is the ordinary least squares solution, ${\bm{w}}^{*}={\bm{\Sigma}}^{-1}{\bm{\beta}}$, which is the global minimum. For linearly separable binary classification with logistic loss, ${\bm{w}}^{*}$ aligns with the max-margin solution, ${\bm{w}}^{*}/\|{\bm{w}}^{*}\|={\bm{w}}_{\mathrm{svm}}/\|{\bm{w}}_{\mathrm{svm}}\|$ where
${\bm{w}}_{\mathrm{svm}}=\underset{{\bm{w}}}{\mathrm{argmin}}\|{\bm{w}}\|^{2}\;\;\mathrm{s.t.}\;\;y_{\mu}{\bm{w}}^{\top}{\bm{x}}_{\mu}\geq 1,\;\forall\;\mu=1,\cdots,P.$ (16)
@@ 4.2 Orthogonal and XOR Datasets
In Section 4.1, we showed that for symmetric datasets satisfying Condition 3, a two-layer bias-free (leaky) ReLU network evolves like a linear network. For more general datasets, the equivalence no longer holds, but comparing ReLU with linear networks remains useful for understanding the learning dynamics of ReLU networks. Specifically, we highlight two cases where a two-layer bias-free ReLU network evolves like multiple independent linear networks. These cases, i.e., an orthogonal input dataset and an XOR-like dataset, are commonly considered in theoretical literature on ReLU network learning dynamics, while their connection to linear network dynamics has not been previously highlighted.
@@
We first look into datasets with orthogonal input, that is $\forall\mu\neq\nu,{\bm{x}}_{\mu}^{\top}{\bm{x}}_{\nu}=0$, a common setting in the literature (Boursier et al., 2022; Telgarsky, 2023; Frei et al., 2023b; c; Kou et al., 2023; Dana et al., 2025). We illustrate with a dataset with two orthogonal data points from Boursier et al. (2022), and handle an arbitrary number of data points in Proposition 18. We train a two-layer bias-free ReLU network on this dataset (LABEL:fig:ortho-dataset), and reproduce the loss curve in (Boursier et al., 2022, Figure 3). We then train two two-layer linear networks on each data point separately. We find that the timing and the amount of the loss drop overlap with the loss curves of the two linear networks as shown in LABEL:fig:ortho-loss. To explain this overlap, we plot the first-layer weights of the ReLU network in black arrows in LABEL:fig:ortho-dataset and find that the weights align with either one of the two data points. Since the two directions are orthogonal, the learning dynamics of the two groups of neurons decouple, as derived in Proposition 18. Each group of neurons evolves like a linear network trained on that single data point. Hence, weights in the ReLU network evolve like a linear network trained on either one of the two data points separately. For a dataset with an arbitrary number of orthogonal inputs, the dynamics of the two-layer bias-free ReLU network evolves like two linear networks trained separately on data points with positive labels and data points with negative labels, as validated in Figure 9. The same applies to learning with logistic loss, as shown in Figure 8.
@@
Figure 4: Two-layer bias-free ReLU networks can evolve like multiple independent linear networks. (a) An orthogonal input dataset used in (Boursier et al., 2022, Figure 3). The $+$ and -- signs represent data points with $+1$ and $-1$ labels respectively. Their different colors are used only to distinguish the loss curves. The black arrows are the first-layer weights at convergence. (b) The loss curve of the ReLU network overlaps with two linear networks trained on each of the two data points respectively. (c) An XOR-like dataset. (d) The loss curve of the ReLU network overlaps with four linear networks trained on each of the four data points separately. Details: We use summed (instead of averaged) square loss for this figure. The initial losses are vertically aligned to help illustrate the overlap. More details are in Appendix H.
@@
We observe similar behaviors in the XOR-like task shown in LABEL:fig:xor-dataset. XOR-like datasets are also a common setting in the literature (Refinetti et al., 2021; Saxe et al., 2022; Frei et al., 2023a; Meng et al., 2024; Xu et al., 2024; Glasgow, 2024). As shown in LABEL:fig:xor-loss, we find that the loss curves of a two-layer bias-free ReLU network trained on the XOR task overlap with four linear networks trained on each data point separately. In this case, the dynamics of multiple linear networks well approximate that of a ReLU network, even though the ReLU network learns a nonlinear function.
@@
In LABEL:fig:ortho-loss and LABEL:fig:xor-loss, the loss curves go through multiple drops, each corresponding to learning a data point. Similar behaviors were examined by Boursier et al. (2022); Xu et al. (2024) and characterized as saddle-to-saddle dynamics. The connections we find between ReLU and linear networks may help understand these behaviors in ReLU networks because saddle-to-saddle dynamics has been well studied for linear networks (Saxe et al., 2014; 2019; Gissin et al., 2020; Jacot et al., 2022; Berthier, 2023; Pesme & Flammarion, 2023).